Given A=[ 1st row; 2nd row]=[-5 2; 2 -2] Which of the following is an eigenvalue of A? -6 6 3 -3
Answer #1


A = -5 2. 2 - 2 Characterstic ean of A 7 | A-AI) = 2 -5-1 2 - 0 2 - 2-1 = (-5-1) (-2-1) -4 = 0 =) (5+12 (2+1) – 4 = 0 10+70+